OpenStudy (anonymous):

Please help! Brad's car travels 12mph faster than Dave's car. in the same time that Brad's car travels 300 miles, Dave travles 240 miles. need the speed of each of their cars. Been racking my brain on this-Please help!!!!!!

OpenStudy (anonymous):

brads car travels 60 mile morw in 1 hr it goes 12 mile more so in 5 hrs it goes 60 mile more so brads spped=300/5=60 and dave =240/5=48